OpenStudy (anonymous):

Simplify the expression. 2(3y - 7) - 5y(2 - y)

OpenStudy (anonymous):

\[5y ^{2}-4y -14\]

jhonyy9 (jhonyy9):

6y-14-10y+5y^2=5y^2 -4y -14

OpenStudy (kira_yamato):

2(3y - 7) - 5y(2 - y) = 2(3y - 7) + 5y(y - 2) = 6y - 14 + 5y^2 - 10y = 5y^2 - 4y - 14